OpenStudy (anonymous):

The surface area of the top surface of the water in a circular swimming pool is about 206 square feet. Estimate the radius of the pool, to the nearest foot.

jhonyy9 (jhonyy9):

so than this mean that area = pi*r^2 = 206 ft^2 so 3,14*r^2 = 206 divide both sides by 3,14 r^2 = 206 / (3,14) = 65,6 than r = sqrt(65,6) = 8,1 so than the radius is approximatly 8 ft